# On the kinematic nature of apparent discs at high redshifts: Local counterparts are not dominated by ordered rotation but by tangentially anisotropic random motion

### v1

@@ 1 Introduction
With statistical analyses of projected shapes of galaxies and their subcomponents (e.g., Padilla & Strauss, 2008; van der Wel et al., 2009; Sánchez-Janssen et al., 2010; Rodríguez & Padilla, 2013; Gao et al., 2020; Xu et al., 2024) and of stellar kinematics (e.g., Emsellem et al., 2011; Krajnović et al., 2011; Cappellari et al., 2013a), now it is generally believed that local galaxy populations dominant in mass are mainly axisymmetric oblate and spheroidal systems, and that the most massive early-type galaxies are mildly triaxial (Weijmans et al., 2014; Li et al., 2018). The fast-rotating oblate and slow-rotating spheroidal/ellipsoidal galaxies manifest a bimodal distribution of their rotational support (Cappellari, 2016; Graham et al., 2018; van de Sande et al., 2021a), which has recently been shown to be a universal property for galaxies across different masses, star formation states, and environments (Wang & Peng, 2023; Wang et al., 2024).
@@ 2 Sample
Milky Way analogs at $z=0$: We also select Milky Way analogs in MaNGA survey to represent the real dynamically cold discs in the local Universe. These 221 analogs are defined by stellar mass range $10.810.6<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 10.8 (around the median of the estimated range in Pillepich et al. (2023)) and are confined within the span of star formation main sequence (SFMS) of local galaxies (section 2.1 of Wang & Peng, 2023). We do not additionally apply any morphology cut as these galaxies are already overwhelmed by fast-rotating discs with high rotational support (Wang et al., 2024).
@@ 5 Discussion and conclusion
Figure 4: The probability density distributions of the ellipticity $\varepsilon$ of galaxy projected shapes, for low-mass galaxies (black solid line) and Milky Way analogs (blue dashed line) observed at $z=0$ and for galaxies in the early Universe $z=3-8$ ($9.59<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 9.5 in JWST-CEERS studied by Pandya et al. (2024); the purple shaded area). The density distributions are derived via kernel density estimation with Gaussian kernels of widths determined based on Scott’s rule. For local low-mass galaxies, we estimate the uncertainty of the distribution using 1000 bootstrap samples shown by the thin black lines. Suggesting the intrinsic shapes to first order, we have also shown the predicted distributions of oblate ($c<b=a,c/a\sim\mathcal{N}(\mu=0.3,\sigma=0.1)$; the thick gray line) and prolate ($c=b<a,c/a\sim\mathcal{N}(\mu=0.3,\sigma=0.1)$; the orange shaded area) ellipsoids randomly projected on the sky plane.
@@
From theoretical point of view, the stellar kinematics of actively star-forming galaxies at higher redshifts are governed more by random motion than their local analogs of equal masses (EAGLE, TNG100, TNG50; Trayford et al., 2019; Tacchella et al., 2019; Pillepich et al., 2019), because the fast accumulated gas becomes gravitationally unstable toward star formation far ahead of complete gas cooling and settling down into a disc (Mo et al., 2023). This is consistent with observations that molecular and ionized gas, which stay closely with young stars, are dynamically hotter at higher redshifts (e.g., Wisnioski et al., 2015). The typical JWST galaxies observed at high $z$ are thus expected to be supported even more by random motion than the local equal-mass counterparts highlighted in this work which were formed recently and quiescently. For the local counterparts, if we focus on those with highest SFRs (0.5 dex above the SFMS; orange circles in Fig. 2), that are more analogous to the high-$z$ population in terms of star formation and gas properties, the median stellar spin does decrease further by 40%.
There are indeed other pieces of evidence for high-$z$ star-forming galaxies being dynamically warm/hot systems with large velocity anisotropy. Most notably, high-$z$ galaxies of relatively low masses feature an asymmetric distribution of projected axis ratio that is biased toward small values. This is the characteristic of intrinsically elongated systems. Zhang et al. (2019) infer the intrinsic shapes of galaxies by modelling the projected axis ratios and sizes based on HST images, and find that in the redshift range $2.52<z<$ 2.5, 70% of $9.59<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 9.5 galaxies and 60% of $109.5<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 10 galaxies are prolate, and the discy population comprise fewer than 20%. The prevalence of low-mass prolate galaxies at high redshifts is confirmed by the latest JWST results (Pandya et al., 2024), showing that about 70% of $109<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 10 galaxies are prolate at $82<z<$ 8. Pandya et al. (2024) also find that high-probability prolate and oblate candidates have remarkably similar Sérsic indices around one and non-parametric morphological properties. With integral field spectroscopy, our work suggests that these prolate systems, unlike rotation-supported discs in oblate shapes, are governed by anisotropic random motion.
@@
So, both having low rotational support, the intrinsic shapes of low-mass galaxies with $\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})\sim 9$ have transitioned from prolate at $z>2$ to oblate at present. Fig. 4 illustrates this transition with the ellipticity distributions of local ($z=0$; black line) and distant ($z=3-8$; purple shaded area; taken from Pandya et al. (2024)) low-mass galaxies. Apparently, the projected shapes of local and distant low-mass galaxies are significantly different, with the distant low-mass galaxies showing much more highly flattened shapes on the sky. The distribution of local low-mass galaxies follows that of the Milky Way analogs (blue dashed line), the real dynamically cold discs, but the distribution of Milky Way analogs is slightly shifted to higher values. Analytical predictions for randomly oriented oblate (the gray line) and prolate (orange shaded area) ellipsoids, both assuming a Gaussian distribution $\mathcal{N}(\mu=0.3,\sigma=0.1)$ for the axial ratio $\xi\equiv c/a$, match the observations in remarkable details. The transition of intrinsic shapes is associated with a change in the type of velocity anisotropy. While the local low-mass galaxies are suggested to possess tangentially biased stellar orbits, the stellar random motions of high-$z$ low-mass galaxies are radially biased, probably along the predominant feeding filament of dark matter as shown in Ceverino et al. (2015).
@@ Acknowledgements
YP and BW acknowledge the National Science Foundation of China (NSFC) Grant No. 12125301, 12192220, 12192222, and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-A07.

### v2

@@ 1 Introduction
With statistical analyses of projected shapes of galaxies and their subcomponents (e.g., Padilla & Strauss, 2008; van der Wel et al., 2009; Sánchez-Janssen et al., 2010; Rodríguez & Padilla, 2013; Gao et al., 2020; Xu et al., 2024) and of stellar kinematics (e.g., Emsellem et al., 2011; Krajnović et al., 2011; Cappellari et al., 2013a), now it is generally believed that local galaxy populations dominant in mass are mainly axisymmetric oblate and spheroidal systems, and that the most massive early-type galaxies are mildly triaxial (Weijmans et al., 2014; Li et al., 2018). The fast-rotating oblate and slow-rotating spheroidal/ellipsoidal galaxies manifest a bimodal distribution of their rotational support (Cappellari, 2016; Graham et al., 2018; van de Sande et al., 2021a), which has recently been shown to be a universal property for galaxies across different masses, star formation states, and environments (Wang & Peng, 2023; Wang et al., 2023).
@@ 2 Sample
Milky Way analogs at $z=0$: We also select Milky Way analogs in MaNGA survey to represent the real dynamically cold discs in the local Universe. These 221 analogs are defined by stellar mass range $10.810.6<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 10.8 (around the median of the estimated range in Pillepich et al. (2023)) and are confined within the span of star formation main sequence (SFMS) of local galaxies (section 2.1 of Wang & Peng, 2023). We do not additionally apply any morphology cut as these galaxies are already overwhelmed by fast-rotating discs with high rotational support (Wang et al., 2023).
@@ 5 Discussion and conclusion
Figure 4: The probability density distributions of the ellipticity $\varepsilon$ of galaxy projected shapes, for low-mass galaxies (black solid line) and Milky Way analogs (blue dashed line) observed at $z=0$ and for galaxies in the early Universe $z=3-8$ ($9.59<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 9.5 in JWST-CEERS studied by Pandya et al. (2024a); the purple shaded area). The density distributions are derived via kernel density estimation with Gaussian kernels of widths determined based on Scott’s rule. For local low-mass galaxies, we estimate the uncertainty of the distribution using 1000 bootstrap samples shown by the thin black lines. Suggesting the intrinsic shapes to first order, we have also shown the predicted distributions of oblate ($c<b=a,c/a\sim\mathcal{N}(\mu=0.3,\sigma=0.1)$; the thick gray line) and prolate ($c=b<a,c/a\sim\mathcal{N}(\mu=0.3,\sigma=0.1)$; the orange shaded area) ellipsoids randomly projected on the sky plane.
@@
From theoretical point of view, the stellar kinematics of actively star-forming galaxies at higher redshifts are governed more by random motion than their local analogs of equal masses (EAGLE, TNG100, TNG50; Trayford et al., 2019; Tacchella et al., 2019; Pillepich et al., 2019), because the fast accumulated gas becomes gravitationally unstable toward star formation far ahead of complete gas cooling and settling down into a disc (Mo et al., 2024). This is consistent with observations that molecular and ionized gas, which stay closely with young stars, are dynamically hotter at higher redshifts (e.g., Wisnioski et al., 2015). Significant rotation of cool and cold gas can still be established at extremely high redshifts (e.g., Fujimoto et al., 2024; Rowland et al., 2024), particularly for massive halos as simulation suggested, but tends to be fragile and destructed after several orbital time (Dekel et al., 2020).
The typical JWST galaxies observed at high $z$ are thus expected to be supported even more by random motion than the local equal-mass counterparts highlighted in this work which were formed recently and quiescently. For the local counterparts, if we focus on those with highest SFRs (0.5 dex above the SFMS; orange circles in Fig. 2), that are more analogous to the high-$z$ population in terms of star formation and gas properties, the median stellar spin does decrease further by 40%.
There are indeed other pieces of evidence for high-$z$ star-forming galaxies being dynamically warm/hot systems with large velocity anisotropy. Most notably, high-$z$ galaxies of relatively low masses feature an asymmetric distribution of projected axis ratio that is biased toward small values. This is the characteristic of intrinsically elongated systems. Zhang et al. (2019) infer the intrinsic shapes of galaxies by modelling the projected axis ratios and sizes based on HST images, and find that in the redshift range $2.52<z<$ 2.5, 70% of $9.59<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 9.5 galaxies and 60% of $109.5<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 10 galaxies are prolate, and the discy population comprise fewer than 20%. The prevalence of low-mass prolate galaxies at high redshifts is confirmed by the latest JWST results (Pandya et al., 2024a), showing that about 70% of $109<\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})<$ 10 galaxies are prolate at $82<z<$ 8. Pandya et al. (2024a) also find that high-probability prolate and oblate candidates have remarkably similar Sérsic indices around one and non-parametric morphological properties. With integral field spectroscopy, our work suggests that these prolate systems, unlike rotation-supported discs in oblate shapes, are governed by anisotropic random motion.
@@
Both having low rotational support, the intrinsic shapes of low-mass galaxies with $\mathrm{lg}(\mathcal{M}_{\star}/\mathcal{M}_{\odot})\sim 9$ have transitioned from prolate at $z>2$ to oblate at present. Fig. 4 illustrates this transition with the ellipticity distributions of local ($z=0$; black line) and distant ($z=3-8$; purple shaded area; taken from Pandya et al. (2024a)) low-mass galaxies. Apparently, the projected shapes of local and distant low-mass galaxies are significantly different, with the distant low-mass galaxies showing much more highly flattened shapes on the sky. The distribution of local low-mass galaxies follows that of the Milky Way analogs (blue dashed line), the real dynamically cold discs, but the distribution of Milky Way analogs is slightly shifted to higher values. Analytical predictions for randomly oriented oblate (the gray line) and prolate (orange shaded area) ellipsoids, both assuming a Gaussian distribution $\mathcal{N}(\mu=0.3,\sigma=0.1)$ for the axial ratio $\xi\equiv c/a$, match the observations in remarkable details. The transition of intrinsic shapes is associated with a change in the type of velocity anisotropy. While the local low-mass galaxies are suggested to possess tangentially biased stellar orbits, the stellar random motions of high-$z$ low-mass galaxies are radially biased, probably along the predominant feeding filament of dark matter (Ceverino et al., 2015; Pandya et al., 2019, 2024b).
@@ Acknowledgements
We appreciate the thoughtful and constructive comments by our reviewer. YP and BW acknowledge the National Science Foundation of China (NSFC) Grant No. 12125301, 12192220, 12192222, and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-A07.